# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v1

@@ 3. CataractBot System Design
3.1. Design Requirements
Drawing from our discussions with doctors and patient coordinators, literature review, and understanding of the Indian healthcare ecosystem, we have identified several key requirements for the chatbot.
Accuracy: The chatbot’s capability to answer both medical and logistical questions with accuracy is a fundamental requirement. Providing precise, contextually relevant, and verified responses is paramount given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it might not be tailored to India-specific or hospital-specific nuances. For instance, while the phacoemulsification method is prevalent globally for cataract surgeries, India often favors manual small incision-based procedures due to their cost-effectiveness (Singh et al., 2017). Similarly, cultural considerations significantly impact advice on pre- and post-surgery choices, e.g., questions related to diet necessitates inclusion and exclusion of Indian food options for accurate guidance. To ensure accurate and contextual response, our chatbot generates responses from a knowledge base curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
Trustworthiness: In medical surgical treatment, access to reliable trustworthy information is crucial. Previous research (Xiao et al., 2023) highlights how patients often encounter misinformation from Internet-based (medical) sources, leading to skepticism towards them. Consequently, they tend to rely on the experiences of friends and family members who have undergone similar treatments, although such resources can sometimes be challenging to locate (Civan-Hartzler et al., 2010). Typically, patients place their utmost trust in the operating doctor and the hospital staff responsible for their care during the procedure (Larasati et al., 2023). Hence, each response the patient (or their attendant) receives is vetted and, if necessary, edited by doctors and patient coordinators to ensure trustworthiness.
Timeliness: Prompt responses are important as delays often lead patients to seek information elsewhere. Patient coordinators revealed that patients frequently reach out to them (over a phone call) for relaying medical queries to their consulting doctors. However, due to busy schedules of doctors and coordinators, immediate responses are not always feasible. That often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation. To counter this, we decided the chatbot should provide instantaneous responses to every patient query using a custom knowledge base curated by hospital staff. Subsequently, these responses were reviewed and refined by doctors/coordinators later.
Usability: The developed system must be usable by patients, their attendants, doctors and patient coordinators. Thus, we chose WhatsApp as our platform to develop CataractBot, as it is the most widely used messaging platform in India (Statista, 2022). Despite Telegram’s richer feature set or a custom chatbot offering more development flexibility, we opted for WhatsApp to minimize the learning curve and ensure seamless onboarding without requiring additional app installations. Furthermore, due to the frequent use of WhatsApp among Indians (Statista, 2022), we anticipated higher adoption and usage of a WhatsApp-based chatbot. To cater to the diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, we made the chatbot multilingual—supporting five languages (English, Hindi, Kannada, Tamil, and Telugu)—and multimodal (offering both text and speech inputs and outputs), similar to (Jain et al., 2018; Findlater et al., 2009).
@@ 3.2. Components
From our discussions with doctors and hospital staff, we have identified four key stakeholders in the cataract surgery ecosystem: the patient, their attendants, the doctor, and the patient coordinator. Once onboarded on CataractBot, a patient (and their attendant) scheduled for cataract surgery within the next 7 days can ask any question related to the surgical treatment and receive a response verified by the hospital staff. In addition, we have a few more roles: an escalation doctor, an escalation patient coordinator, and a KB (Knowledge Base) update expert. Below, we describe the key components of our CataractBot system.
@@ 3.2.1. Input Language and Modality
The Sankara Eye Hospital situated in Bangalore (in the state of Karnataka), known as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. To accommodate this diverse patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language via WhatsApp. Among these messages is an option allowing users to modify their preferred language using a drop-down menu at any time in the future. Note: To converse in a specific language, users must ensure their preferred language is selected before interacting. For example, if a user has chosen English as their preferred language, they cannot begin conversing with CataractBot in Kannada. However, they can navigate to the welcome messages, switch their preferred language to Kannada, and then engage with the bot in Kannada.
To accommodate different levels of literacy, including literate, semi-literate, and illiterate users, both text and speech inputs are enabled (Jain et al., 2018; Findlater et al., 2009). Users can ask questions by holding the mic button on WhatsApp, speaking in their preferred language, and sending a voice message. For every voice message, CataractBot not only provides a text response but also sends an audio message (Figure 2D). Given that the doctor and patient coordinator may not be proficient in all five languages, their interactions with the bot for verifying and editing messages are exclusively conducted in English.
Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot addresses this by providing a set of three frequently asked questions in a drop-down menu along with the welcome message. Additionally, after every medical and logistical question from the patient/attendant, the bot’s response includes a “What to do next?” prompt, offering three related questions in a drop-down menu based on the preceding query (Figure 2F). This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves the user time. Users are free to select one of the suggested questions or pose a different query to continue the conversation.
@@ 3.2.2. Response Generation, Verification and Icons
Upon receiving a message from a patient/attendant, CataractBot classifies it as a medical question, logistical question, small talk, or other queries.
For medical and logistical questions, the bot strictly employs the knowledge base curated by the hospital’s medical team to generate an appropriate response.
In instances where the custom knowledge base lacks an answer, the bot responds with a template “I don’t know” response.
@@
For miscellaneous questions unrelated to medical or logistical matters, e.g., about sports or weather, the bot responds with a template “I don’t know” message.
The patient/attendant receives these responses in real-time from CataractBot.
Specifically, for medical and logistical questions, the response includes a question mark icon
as a reaction indicating the unverified status (Figure 2E).
For medical questions, the operating doctor of the respective patient receives a message comprising of the question asked, the bot’s response, patient demographics, and citations of the documents used to generate the response.
The doctor is then prompted with the question, “Is the answer accurate and complete?” offering three response options: Yes, No, and Send to Patient Coordinator (Figure 1A).
@@
in the user’s received answer with a green tick icon
, confirming verification (Figure 1F). Additionally, the bot notifies the user that the answer has been verified, tagging that particular response (Figure 1G).
@@
, indicating an incorrect answer (Figure 1C), and asks the user to await a corrected response from the doctor (Figure 1D).
The doctor is asked to provide a correction; they are not required to edit the bot’s message, instead, they need to offer a correction in free form.
@@
Similar to the doctor’s workflow, for logistical questions, the patient coordinators verify and provide corrections to the bot’s response.
In deciding how to deliver chatbot responses to patients, we deliberated between displaying unverified responses in real-time and presenting only verified responses post-expert verification and correction. Through our discussions with hospital staff, it became clear that providing real-time responses was crucial. Delayed responses might lead to patients making repeated calls to the hospital or resorting to online sources, failing to effectively address their information needs.
@@ 3.2.3. Escalation Mechanism
The CataractBot system employs an intricate escalation mechanism to ensure swift verification of every query posed by the patient/attendant.
@@
(Figure 2E), enabling them to easily identify pending queries.
If an answer is not verified by the operating doctor within three hours, it automatically escalates to the designated escalation doctor.
This duration was set based on input from the hospital staff.
@@
Selecting ‘Yes’ or ‘Send to Patient Coordinator’ immediately marks the query with a green tick
(Figure 1F) for both the doctors, indicating task completion.
@@
In instances where neither the operating nor escalation doctor verifies a question within six hours, both receive a notification indicating the pending status.
Additionally, every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours is sent to both the operating and escalation doctors, tagged with appropriate messages.
This additional reminder was added based on the doctor’s feedback.
This workflow is mirrored for the patient coordinator and escalation coordinator.
@@ 3.2.4. KB Update Process
While verification remains vital in the medical context, we aim to reduce the effort required from already burdened medical professionals in verifying and editing CataractBot responses. To achieve this, we devised a mechanism to minimize repetitive edits performed by experts. This process involves using the expert-provided edits to update the KB, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future. However, we understand that certain responses, such as those specific to individual patients (e.g., “when to reach the hospital for my surgery?”), should not be added to the KB. Thus, we appointed a senior cataract surgery doctor as the ‘KB update expert’. Each night at 8 pm, this expert automatically receives an online spreadsheet via email containing columns such as ‘Question asked’, ‘Bot’s response’, ‘Expert’s correction’, ‘Bot’s Corrected Response’, and ‘Should Update KB?’. Their task involves reviewing each row and determining whether the information should be added to the CataractBot KB. They respond with a ‘Yes’ or ‘No’ in the ‘Should Update KB?’ column. If marked ‘Yes’, the KB update expert can modify the ‘Bot’s Corrected Response’ column as needed. At 3 am daily, we extract ‘Question asked’ and ‘Bot’s Corrected Response’ data from all rows marked ‘Yes’, and append these to the KB in the ‘expert-FAQ’ document. Additionally, we prompt the large language model (LLM) to prioritize this ‘expert-FAQ’ document within the KB. This ensures that the bot learns from these updates, gradually enhancing its knowledge over time. Consequently, when the bot encounters a similar question in the future, it can provide an accurate response without requiring further corrections, thus effectively reducing the workload of doctors and coordinators.
@@ 3.3. Implementation Details
The CataractBot system mainly relies on these five services (Figure 3): (a) LLM (Large Language Model): for response generation, (b) Vector Database: for storing and retrieving the custom knowledge base, (c) Language Technologies: for translation, and for converting speech-to-text and text-to-speech, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
@@
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). Considering the medical context, instead of relying on GPT-4’s default knowledge, CataractBot utilized a curated knowledge base by the hospital staff. This custom KB comprises of various Cataract surgery and Sankara Eye Hospital-related documents, including Consent Form, SOP (Standard Operating Procedure), FAQs (Frequently Asked Questions), Pre- and Post-Operative Guidelines, Hospital Information, and Patient Discharge Do’s and Dont’s. These documents are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question asked by patient/attendant, GPT-4 classifies the query type (medical, logistical, small talk, or other). For a medical/logistical question, CataractBot employs a RAG (Retrieval Augmented Generation) (Lewis et al., 2020) approach. This involves performing a vector search on the KB to extract the three most relevant data chunks related to the query. GPT-4 is then prompted (full prompt available in Appendix A) to extract an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message. Moreover, GPT-4 is prompted to give priority to the ‘expert-FAQ’ document–vetted by medical professionals–when it is among the top data chunks. The prompt also instructs that the response should adhere to WhatsApp’s message limit of 700 characters.
@@
This prompt underwent several iterations. We engaged a doctor and a patient coordinator at the hospital to record questions posed by patients/attendants scheduled for cataract surgery over a week. A total of 153 questions were compiled. We used our GPT-4 prompt to obtain answers to these questions from the custom KB, and subsequently the generated responses were evaluated by the same doctor and coordinator. This not only helped refine the LLM prompt, but also led the medical staff to improve the content of existing documents and contribute additional documents to enrich the KB. This process was iterated three times, to ensure comprehensive coverage and validate the effectiveness of our GPT-4 prompt with the RAG approach in extracting pertinent responses. Despite their effectiveness, we recognized the KB’s susceptibility to outdated or incomplete information. Hence, expert edits to bot’s answers, confirmed by the KB update expert, were consistently incorporated to regularly enhance and update the KB.
Additionally, we utilized GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 flawlessly executed this task.
Discussing language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, the CataractBot system adhered to the standard approach (ait, [n. d.]) of translating local Indic languages (Hindi, Kannada, Tamil, and Telugu) into English for input using Azure AI Translator. It then translated GPT-4’s English responses back into the respective local Indic languages for output. Moreover, to facilitate speech as input and audio as output, CataractBot used Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.)
During testing, we encountered instances where GPT-4 was busy, resulting in response failures due to timeouts. To guarantee responses for every question, we implemented a queuing system. Each message from patients/attendants entered a queue and was removed only after receiving a response from GPT-4. Furthermore, to ensure expert verification, we employed multiple CronJobs for notifications and escalations. Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and KB update expert) gets logged in a cloud storage for further analysis.
@@ 4. Study Design
To evaluate the effectiveness of an LLM-powered expert-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 in collaboration with a local eye hospital, Sankara Eye Hospital. Sankara Eye Hospital is one of the leading tertiary eye care and teaching institutions in Bengaluru, India. It has 18 eye doctors (15 of whom perform cataract surgeries regularly) and over 10 optometrists. The hospital typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital.
@@ 4.1. Procedure for Patients and Attendants
As per the hospital’s protocol, once a patient elects to proceed with cataract surgery following the doctor’s recommendation, the patient and their attendant are directed by the doctor to meet a patient coordinator. The coordinator helps with general counseling, guidance on pre- and post-operative measures, scheduling of the surgery date, and logistical arrangements. Toward the end of this interaction, the patient coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages (English, Hindi, Kannada, Tamil, or Telugu), (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as they would likely have minimal informational needs), and (e) being a patient of one of the three participating operating doctors. If the patient meets these criteria, the patient and their attendant are directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
@@
The researcher introduced them to CataractBot and outlined the study’s protocol which involved using CataractBot for around two weeks, engaging in two interviews, and completing a chatbot usability form. Upon confirming their willingness to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form. This form included details such as the patient and attendant’s phone numbers linked to WhatsApp, preferred languages, consulting doctor and coordinator, surgery date, and basic demographic information (age, gender, and education). Upon form submission, participants received a set of ‘welcome messages’ from CataractBot. The researcher instructed them to ask a trial question to the bot, either by choosing from suggested questions (Figure 2F) or by typing/speaking a new question in their preferred language. Upon receiving a response, participants were briefed on the three icons (Figure 1E,F): question mark, green tick, and red cross. Throughout this process, the researcher encouraged them to clarify any doubts regarding the bot’s usage. Following this onboarding process, patients and their attendants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to ask questions to CataractBot.
@@
Participants were instructed to use the bot from onboarding to one week post-surgery, after which their access was automatically revoked. Given that cataract surgeries were typically scheduled within a week of meeting with the patient coordinator, each patient/attendant was expected to use the bot for a maximum of two weeks. We conducted two interviews with both the patient and their attendant: one on the surgery day and another a week post-surgery. These semi-structured interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses. Both interviews followed a similar structure, differing only in how the bot supported the participant before versus after surgery. These specific days—surgery day and seven days post-surgery—were chosen because they coincided with the patient’s required hospital visits. After each interview, participants were requested to fill out a chatbot usability form. They rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness. These adapted questions were sourced from metrics outlined in prior usability evaluations of healthcare bots (Abd-Alrazaq et al., 2020). Patient-attendant pairs were interviewed together but were asked to fill separate usability forms.
@@
All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages. (Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while the exit interviews (seven days post-surgery) were conducted either in person at the hospital or through pre-scheduled telephone calls. This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with surgery preparations and the procedure itself taking 45 minutes. Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time. Each interview typically spanned 20 to 40 minutes and was audio-recorded with the participant’s consent.

### v2

@@ 4. CataractBot System Design
Here, we describe the CataractBot system (Figure 3), including its key components–input language and modality, response generation and verification, and escalation and reminders–and provide a detailed account of its implementation.
@@ 4.1.1. Input Language and Modality
Sankara Eye Hospital in Bangalore (in the state of Karnataka), referred to as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. Analysis of a 2011 census highlights Bangalore as one of India’s most linguistically diverse cities (Office of the Registrar General & Census Commissioner, 2011). To accommodate this diverse patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). As the expert may not be proficient in all five languages, their interactions with the bot are exclusively in English. Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language. Among these messages is an option allowing users to modify their preferred language in future. To accommodate different levels of literacy, including literate, semi-literate, and illiterate users, CataractBot supports both text and speech inputs. For every voice message, CataractBot sends both a text response and a spoken version of the text as an audio message.
Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot provides a set of three frequently asked questions along with the welcome message. Additionally, every bot’s response includes a “What to do next?” prompt, offering three related questions based on the preceding query (Figure 2F). This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves the user time.
@@ 4.1.2. Response Generation, Verification and Icons
Upon receiving a message, CataractBot classifies it as a medical question (e.g., dos and donts before and after surgery), logistical question (e.g., scheduling or insurance related), small talk, or ‘other’, and provides a response in real-time.
For medical and logistical questions, the bot strictly employs the KB curated by the hospital’s medical team to generate an appropriate response.
The response includes a question mark icon
as a ‘reaction’ indicating the unverified status (Figure 2E).
In instances where the custom KB lacks an answer, the bot responds with a template “I don’t know” response.
@@
For medical questions, the operating doctor of the respective patient receives a message comprising of the question asked, the bot’s response, and patient demographics (Figure 1A).
As Rajashekar et al. (2024) found that citations improve trust in LLM-generated responses, this message also includes citations of the documents used to generate the response.
The doctor is prompted with the question, “Is the answer accurate and complete?” offering three response options: Yes, No, and Send to Patient Coordinator (Figure 1A).
@@
in the information seeker’s received answer with a green tick icon
, confirming verification (Figure 1F). Additionally, the bot notifies the information seeker that the answer has been verified, tagging that particular response (Figure 1G).
@@
, indicating an incorrect answer (Figure 1C), and asks the information seeker to await a corrected response from the doctor (Figure 1D).
The doctor is asked to provide a correction; they are not required to edit the bot’s message, instead, they can offer a correction in free form text.
@@
This new response is not sent to the doctor, to minimize their workload by limiting verification to a single round.
@@
Similar to the doctor’s workflow, the patient coordinators verify and provide corrections to the bot’s response for logistical questions.
We deliberated between displaying unverified responses in real-time and presenting only verified responses post-expert verification and correction.
Through our
formative study and discussions with hospital staff, it became clear that providing real-time responses was crucial.
Delayed responses might lead to patients making repeated calls to the hospital or resorting to online sources, failing to effectively address their information needs.
@@ 4.1.3. Escalation Mechanism
The CataractBot system employs an intricate escalation and reminder mechanism to ensure swift verification.
@@
(Figure 2E) and answered questions with a green tick
, enabling them to easily identify pending queries.
If an answer is not verified by the operating doctor within three hours, it is automatically sent to the designated escalation doctor.
@@
Selecting ‘Yes’ or ‘Send to Patient Coordinator’ immediately marks the query with a green tick (Figure 1F) for both the doctors, indicating task completion.
@@
If neither the operating nor escalation doctor verifies a question within six hours, both receive a reminder notification indicating the pending status.
Additionally, every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours is sent to both the operating and escalation doctors.
This additional reminder was added based on the doctors’ feedback.
This workflow is mirrored for the patient coordinator and escalation patient coordinator.
@@ 4.1.4. KB Update Process
To minimize experts’ labor, we use the expert-provided edits to update the KB, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future. However, certain responses, such as those specific to individual patients (e.g., “When to reach the hospital for my surgery?”) must not be added to the KB as they are not generalizable. We appointed a senior cataract surgeon as the ‘KB update expert’, who received a spreadsheet via email at 8pm daily. Their task involves reviewing each row and determining whether the information should be added to the KB (by responding with a ‘Yes’/‘No’ in the ‘Should Update KB?’ column), and if so, modifying as needed the ‘Final Answer for KB’ column, containing the bot’s updated answer based on the expert’s correction. At 3 am daily, the system extracts ‘Question’ and ‘Final Answer for KB’ data from all rows marked ‘Yes’, and append these to the KB in the ‘expert-FAQ’ document. Additionally, we prompt the LLM to prioritize this ‘expert-FAQ’ document within the KB. This ensures that CataractBot improves with these updates, gradually enhancing its accuracy over time.
@@ 4.2. Implementation Details
The CataractBot system relies on these five components (Figure 3): (a) LLM: for response generation, (b) Vector Database: for storing and retrieving the custom KB, (c) Language Technologies: for translation and transcription, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
@@
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). The custom KB comprises of various cataract surgery and hospital-specific documents, including Consent Form, SOP (Standard Operating Procedure), FAQs, Pre- and Post-Operative Guidelines, and Hospital Information. These documents are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question, GPT-4 classifies the query type. For a medical/logistical question, CataractBot employs a RAG (Lewis et al., 2020) approach. This involves performing a vector search on the KB to extract the three most relevant data chunks related to the query. GPT-4 is then prompted (full prompt available in Appendix A) to generate an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message.
@@
The prompt underwent several iterations. To validate its effectiveness, a doctor and a patient coordinatorrecorded 153 questions posed by patients/attendants scheduled for cataract surgery over a week. We used our GPT-4 prompt to obtain answers to these questions from the custom KB, and subsequently the generated responses were evaluated by the same doctor and coordinator. This not only helped refine the LLM prompt, but also led the medical staff to improve existing documents and contribute additional documents to enrich the KB. This process was iterated thrice.
Additionally, we utilize GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 executed this task well. Regarding language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, our system adhered to the standard approach (ait, [n. d.]) of translating Indic languages into English for input using Azure AI Translator. It then translated GPT-4’s English responses back into the respective Indic languages for output. To facilitate speech as input and audio as output, CataractBot uses Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.) Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and KB update expert) get logged in a cloud storage for further analysis.
@@ 5. Study Design
To evaluate the effectiveness of this LLM-powered experts-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 at our collaborators’ institute, Sankara Eye Hospital. This is one of the leading tertiary eye care and teaching institutions in Bangalore, India. It typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital. None of the study participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation, in accordance with hospital norms.
@@ 5.1. Procedure for Patients and Attendants
As per the hospital’s protocol, once a patient opts for cataract surgery based on their doctor’s recommendation, the patient and their attendant needs to meet a patient coordinator. The coordinator helps with general guidance on pre- and post-operative measures, scheduling of the surgery date, and logistical arrangements. Toward the end of this interaction, the coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages supported by CataractBot, (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as recent patients would likely have minimal informational needs), and (e) having one of the three participating operating doctors as their surgeon. If these criteria were met, the patient and their attendant were directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
@@
The researcher introduced them to CataractBot and outlined the study’s protocol which involved using the bot for around two weeks, engaging in an interview pre and post-surgery, and completing a chatbot usability form. If they agreed to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form. This form included details such as the patient and attendant’s phone numbers linked to WhatsApp, preferred languages, consulting doctor and coordinator, surgery date, and basic demographic information (age, gender, and education). Upon form submission, participants received ‘welcome messages’ from CataractBot. The researcher instructed them to ask a trial question, either by choosing from suggested questions or by typing/speaking in their preferred language. Upon receiving a response, participants were briefed on the icons and expert verification system. Throughout this process, the researcher encouraged them to ask any questions regarding the bot’s usage. After onboarding, participants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to “Feel free to ask any cataract surgery related questions” until one week post surgery, when their access was automatically revoked.
@@
We conducted two semi-structured interviews with both the patient and their attendant: one on the surgery day and another a week post-surgery. These specific days were chosen because they coincided with the patient’s required hospital visits. The interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses. Both interviews followed a similar structure, differing only in how the bot supported the participant before versus after surgery. After each interview, participants were requested to fill out a chatbot usability form. They rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness (Appendix A.2). These questions were adapted from metrics outlined in prior usability evaluations of healthcare bots; previous work used scales that included aspects not relevant to our study, such as the chatbot’s visual design (Abd-Alrazaq et al., 2020; Holmes et al., 2023). Patient-attendant pairs were interviewed together but were asked to fill separate usability forms.
@@
All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages. (Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while seven days post-surgery interviews were conducted either in person at the hospital or through pre-scheduled telephone calls. This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with the procedure itself taking 45 minutes. Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time. Each interview typically spanned 20 to 40 minutes and was audio-recorded with the participant’s consent.